# VEnvision3D: A Synthetic Perception Dataset for 3D Multi-Task Model Research

### 29 Feb 2024 v1

@@ 4 Baseline Experiments
In this section, we tested multiple subtasks provided by VEnvision3D using the current mainstream methods. For fair comparison, we faithfully follow the experimental settings of each baseline in the original publication.
Depth completion. We carried out experiments on the projection maps of different channal numbers using deep learning methods [45, 24] and traditional methods [74]. Tab. 4 presents our experimental results. PENet achieved the best results, while the other two methods performed poorly. However, it should be noted that our ground truth was generated by a renderer and is denser than the annotated depth maps in the KITTI DC dataset, so the accuracy of these methods on our dataset is much lower than the results on the KITTI DC dataset. Additionally, we observed that as the LiDAR channel number increased, the accuracy of all methods improved to varying degrees, which was particularly evident in NLSPN. This revealed the limitations of existing methods and demonstrated the enhancing effect of increasing point density on depth completion tasks. Fig. 3 shows the results of PENet for data with different point densities. It can be observed from both quantitative and visual results that with the increase in point density, there is a slight improvement in the accuracy of the completion results.
Method ${}^{{\dagger}}$ RMSE MAE iRMSE iMAE
NLSPN [45] 32 6972.50 4482.53 17.61 13.47
64 6538.66 4434.13 19.27 14.50
128 6021.36 3714.02 16.89 12.31
KITTI 741.68 199.59 1.99 0.84
PENet [24] 32 2833.97 803.90 7.10 2.52
64 2808.22 782.92 6.44 2.81
128 2724.75 739.439 9.21 2.40
KITTI 730.08 210.55 2.17 0.94
Zhao et al. [74] 32 8683.49 5126.70 23.20 16.46
64 8405.77 5001.68 23.01 16.35
128 8283.88 4950.97 23.06 16.36
KITTI 1239.84 298.30 3.76 1.21
Table 4: Benchmarking results of monocular depth completion on the test set. We adopt four metrics for performance evaluation, which are root mean squared error (RMSE [mm]), mean absolute error (MAE [mm]), root mean squared error of the inverse depth (iRMSE [1/km]), and mean absolute error of the inverse depth (iMAE [1/km]). ${}^{{\dagger}}$e conducted experiments using data of varying densities, including 32/64/128-channel LiDAR point cloud projection map.
@@
Figure 3: Depth completion results of PENet from different input(32/64/128-channal LiDAR projected image).
Point cloud place recognition. Our place recognition dataset used the same data preprocessing procedure as benchmark datasets proposed in  [59] (except for removing ground plane points). To investigate the impact of dynamic environments, we generated four sequences before and after simulating environmental changes, denoted as Org. and Mig. respectively. Each run was split into two disjoint reference maps used for training and testing, and each reference map was organized as a set of submaps at regular intervals. We chose PointNetVLAD [59], PPTNet [28] and MinkLoc3dv2 [32] as global feature extractors for place recognition and trained each extractor with 2 different training sets of submaps from Org. only and Org.+Mig. Regarding validation, we uniformly employed cross-validation on all eight sequences. Tab. 5 demonstrated that training PPTNet [28] and MinkLoc3dv2 [32] with two sets of sequence data led to noticeable improvement compared to using only the original sequence. This finding highlights the influence of environmental changes on the accuracy of place recognition for the first time. The provided datasets can support ongoing research on change-robust or change-aware place recognition techniques. However, concurrent use of both training sets for PointNetVLAD resulted in a decrease in performance. We conjecture that this is due to the smaller parameter size of the network.
Method ${}^{{\dagger}}$ AR@1 AR@5
PointNetVLAD [59] Org. 87.7 90.4
Org.+Mig. 84.3 90.3
PPTNet [28] Org. 92.6 94.4
Org.+Mig. 94.1 95.0
MinkLoc3dv2 [32] Org. 89.9 94.9
Org.+Mig. 97.1 98.5
LiDAR odometry estimation. We conducted experiments on three sequences according to the settings of KITTI Odometry. Tab. 6 presents our experimental results, where KISS-ICP [63] achieved the best results in Town05 and Town10, while A-LOAM [71] achieved the best results in Town03. However, it is unusual that MULLS [44] obtained significant deviations in Town03 and Town05, and A-LOAM’s results in Town05 were also unsatisfactory. We also tested ORB-SLAM2 [41] using stereo images, but the final performance was poor, as the algorithm lost track early in the sequence. We speculate that the reason for these low accuracy phenomena is due to the sharp turning angle in the trajectories. Overall, KISS-ICP exhibits the strongest robustness, which aligns with its characteristic of not requiring parameter tuning.
Method Sequence ATE ARE
@@
Scene-level point cloud upsampling. We conducted upsampling experiments using point-based [46] and transformer-based [39] methods. The results are presented in Tab. 7. It is important to mention that including the HD metric in the evaluation metrics does not effectively reveal the superiority or inferiority of the algorithm in scene-level upsampling experiments since HD only reflects the maximum mismatch between point sets. As can be seen, PU-GCN performs better in 16/32-channal input while PC2-PU is better in 64-channal input. As the density of input points increases, the accuracy of prediction results also improves. This characteristic is more prominent in PC2-PU. The specific visual results are shown in Fig. 4. For the first time, we have completed the evaluation of the upsampling task on scene-level point clouds. However, from this result, it can also be seen that compared to traditional 3D object upsampling tasks, the upsampling effect at the scene level is not satisfactory. The details of some large buildings are severely lost, which poses new challenges for the point cloud upsampling task.
Method ${}^{{\dagger}}$ CD HD
@@
Figure 4: Upsampling results of PU-GCN from different input(16/32/64-channal LiDAR scan). The ground truth is 128-channal LiDAR scan.
Method Data mIoU
@@
Point cloud semantic segmentaion. We utilized pointwise-MLP-based [25], point-convolution-based [57] and voxel-based [11] methods to test our dataset. The results are shown in Tab. 8. It can be seen that MinkowskiNet achieved the best results(voxel size is all set to 0.05). As for the pointwise-MLP-based RandLA-Net, the test results for different channel numbers are not satisfactory. For example, for small categories like traffic signs, there are even cases where the IoU is 0. We speculate that this is because the downsampling strategy used in RandLA-Net is not suitable for segmenting small objects in large scenes, resulting in performance degradation. In point-convolution-based KPConv, compared to the 64-channel data, the 128-channel data did not show a significant improvement. This is because, for the KPConv, with higher-density input point clouds, precise adjustment of the neighborhood radius is essential to prevent performance degradation. The multi-density point clouds we offer facilitate robustness testing of semantic segmentation algorithms across domains, aiding researchers in exploring domain adaptation or generalization studies.
@@ 5 Multi-task Exploration
In this section, we take upsampling and segmentation as examples to investigate the mutual reinforcement between different tasks, demonstrating the crucial role of multi-task learning in 3D scene understanding. Based on the results, we discuss the potential of our dataset in establishing end-to-end foundation models within the 3D vision field.
@@ 5.1 The Network
We adopt PU-Dense [2] network architecture as the backbone network and extend a branch for semantic segmentation task. The detailed network architecture is illustrated in Fig. 5. Apart from the original PU-Dense structure, we incorporate transposed convolution [11] to upsample features from different scales of the encoder. These features are then concatenated as input, and the SENet [22] is applied to learn the importance of each channel. Ultimately, a linear classifier is employed to accomplish the semantic segmentation. Besides, we adjusted the expansion factor (kernel size) for coordinate generation within the PU-Dense network to accommodate large-scale scenes.
@@ 5.3 Experiments
To explore the relationships between multiple subtasks, we adopt the following experimental strategies: training solely with the upsampling branch; training solely with the semantic segmentation branch; and simultaneous training with both the upsampling and semantic segmentation branches. Such a design allows us to quantify the phenomena of inter-task influence, providing a research basis for the design of end-to-end foundation models.
Using data with different point densities from our dataset, we take the coordinate values of the 32-channal point cloud as input, the 128-channal point cloud as the ground truth for upsampling, and the label values of the 32-channal point cloud as the ground truth for semantic segmentation. For a fair comparison, the Adam [30] and SGD [52] optimizer is used with a weight decay of 0.0001 and a default learning rate of 0.002. We use exponentialLR [36] to adjust learning rate and the gamma is 0.5. During the training procedure, random rotation along z-axis is performed for the augmentation. The model is trained for 5 epochs with a batch size of 4 under each setting.
Branch Ups. AR Ups. CD Seg. mIoU
@@
The experimental results are shown in Tab. 9. From the experimental results, it can be observed that when the upsampling and segmentation branches are conducted simultaneously, both the Average Recall (AR) of the upsampling task and the mean Intersection over Union (mIoU) of the segmentation task show improvement. They increased from 43.2% to 46.5% and from 43.6% to 45.3%, respectively. Such results demonstrate the phenomenon of mutual promotion among specific downstream subtasks. The upsampling task forces the encoder to capture the implicit surface structure of objects, and such information can provide geometric constraints for semantic segmentation. Conversely, the semantic information provided by the semantic segmentation task can also assist the upsampling task in obtaining category-related information, thereby improving the upsampling process. It is worth noting that the Chamfer Distance (CD) does not meet the expected performance as PU-Dense is a voxel-based method that achieves upsampling by expanding the voxels occupied by the original spatial points. So it does not adequately reflect the quality of the upsampling.
@@
Through simple experimental setups, we can leverage the proposed dataset for multi-task learning. This demonstrates the advantage of collecting the dataset from the same scenarios, allowing for the convenient integration of multiple tasks without the need for complex mappings. This also reflects that our dataset can support the exploration of an end-to-end multi-task framework.
6 Conclusion
@@
We introduce VEnvision3D, a synthetic perception dataset featuring the exploration of 3D multi-task models, including depth completion, segmentation, upsample, place recognition, and 3D reconstruction. In a virtual environment, we have implemented innovative sensor configurations, endowing our dataset with unique characteristics. For instance, it can provide data with varying densities within the same frame, offer surface-sampled data at a city level, and simulate environmental changes. These features aim to contribute to the advancement of the field of point cloud perception. In addition, benefiting from the carefully designed settings in our data collection, we have integrated the upsampling task and semantic segmentation task into a multi-task framework. Through experiments, we have demonstrated the mutually beneficial relationship among different subtasks. In summary, by introducing this dataset, we hope to provide researchers with a platform to study unified multi-task models, thereby advancing the development of foundation models in the 3D vision field.
Supplementary Material
(a) Town01
(b) Town02
(c) Town03
(d) Town04
(e) Town05
(f) Town06
(g) Town07
(h) Town10
Figure 6: The 8 city maps used during the collection process.
(a) 32-channal pred
(b) 64-channal pred
(c) 128-channal pred
(d) 32-channal GT
(e) 64-channal GT
(f) 128-channal GT
Figure 7: Semantic segmentation results of Minkowski U-Net from different input(32/64/128-channal LiDAR scan).
Figure 8: Visualization results of the designed 4 routes in point cloud place recognition task.
(a) A-LOAM
(b) KISS-ICP
(c) MULLS
Figure 9: Visulization results of LiDAR odometry estimation on each sequence. Note that 00 refers to sequence Town03, 01 refers to sequence Town05 and 02 refers to sequence Town10.

### 05 Mar 2024 v2

@@ 4 Baseline Experiments
In this section, we test five sub-tasks provided by VEnvision3D using the current mainstream methods. For fair comparison, we faithfully follow the experimental settings of each baseline in the original publication. From the experimental result, we found the positive impact of varying point cloud densities on semantic segmentation and depth completion tasks, as well as the limitations of current upsampling algorithms for scene-level tasks, and the effect of environmental migration on the robustness of the network in the place recognition task. These findings empirically validate some past conjectures and open new perspectives for developing algorithms with enhanced domain adaptability. The visualization results of some experiments can be found in the supplementary materials.
@@
Figure 4: Depth completion results of PENet from different input (32/64/128-channel LiDAR projected image).
Depth completion. We carry out experiments on the projected image of different channel count using deep learning methods [42, 24] and traditional methods [70]. Tab. 5 presents our experimental results. PENet achieve the best results, while the other two methods performed poorly. However, it should be noted that our ground truth was generated by a renderer and is denser than the annotated depth maps in the KITTI DC dataset, so the accuracy of these methods on our dataset is much lower than the results on the KITTI DC dataset. Additionally, we observed that as the LiDAR channel count increased, the accuracy of all methods improve to varying degrees, which was particularly evident in NLSPN. This reveals the limitations of existing methods and demonstrate the enhancing effect of increasing point density on depth completion tasks. Fig. 4 shows the results of PENet for data with different point densities. It can be observed from both quantitative and visual results that with the increase in point density, there is a slight improvement in the accuracy of the completion results.
Table 4: Benchmarking results of depth completion on the test set. The accuracy are presented as root mean squared error (RMSE [mm]) and mean absolute error (MAE [mm]) ${}^{{\dagger}}$e conducted experiments using data of 32/64/128-channel LiDAR point cloud projected image.
Method ${}^{{\dagger}}$ RMSE normal-$\downarrow$ MAE normal-$\downarrow$
NLSPN [42] 32 6972.50 4482.53
64 6538.66 4434.13
128 6021.36 3714.02
KITTI 741.68 199.59
PENet [24] 32 2833.97 803.90
64 2808.22 782.92
128 2724.75 739.43
KITTI 730.08 210.55
Physical Surface Mod [70] 32 8683.49 5126.70
64 8405.77 5001.68
128 8283.88 4950.97
KITTI 1239.84 298.30
Method Sequence ATE normal-$\downarrow$ ARE normal-$\downarrow$
@@
Table 4: Benchmarking results of depth completion on the test set. The accuracy are presented as root mean squared error (RMSE [mm]) and mean absolute error (MAE [mm]) ${}^{{\dagger}}$e conducted experiments using data of 32/64/128-channel LiDAR point cloud projected image.
@@
LiDAR odometry estimation. We conduct experiments on three sequences according to the settings of KITTI Odometry. Tab. 5 presents our experimental results, where KISS-ICP [59] achieve the best results in Town05 and Town10, while A-LOAM [67] achieve the best results in Town03. However, it is unusual that MULLS [41] obtain significant deviations in Town03 and Town05, and A-LOAM’s results in Town05 were also unsatisfactory. We also test ORB-SLAM2 [40] using stereo images, but the final performance was poor, as the algorithm lost track early in the sequence. We speculate that the reason for these low accuracy phenomena is due to the sharp turning angle in the trajectories. Overall, KISS-ICP exhibits the strongest robustness, which aligns with its characteristic of not requiring parameter tuning.
Point cloud place recognition. Our place recognition dataset use the same data preprocessing procedure as benchmark datasets proposed in [55] (except for removing ground plane points). To investigate the impact of dynamic environments, we generate four sequences before and after simulating environmental changes, denoted as Org. and Mig. respectively. Each run is split into two disjoint reference maps used for training and testing, and each reference map is organized as a set of submaps at regular intervals. We choose PointNetVLAD [55], PPTNet [28] and MinkLoc3dv2 [32] as global feature extractors for place recognition and train each extractor with 2 different training sets of submaps from Org. only and Org.+Mig. Regarding validation, we uniformly employ cross-validation on all eight sequences. Tab. 7 demonstrated that training PPTNet [28] and MinkLoc3dv2 [32] with two sets of sequence data led to noticeable improvement compared to using only the original sequence. This finding highlights the influence of environmental changes on the accuracy of place recognition for the first time. VEnvision3D datasets can support ongoing research on change-robust or change-aware place recognition techniques. However, concurrent use of both training sets for PointNetVLAD results in a decrease in performance. We conjecture that this is due to the smaller parameter size of the network.
Method ${}^{{\dagger}}$ AR@1 normal-$\uparrow$ AR@5 normal-$\uparrow$
PointNetVLAD [55] Org. 87.7 90.4
Org.+Mig. 84.3 90.3
PPTNet [28] Org. 92.6 94.4
Org.+Mig. 94.1 95.0
MinkLoc3dv2 [32] Org. 89.9 94.9
Org.+Mig. 97.1 98.5
Method ${}^{{\dagger}}$ CD normal-$\downarrow$ HD normal-$\downarrow$
@@
Figure 5: Upsampling results of PU-GCN from different input (16/32/64-channel LiDAR scan). The ground truth is 128-channel LiDAR scan.
Scene-level point cloud upsampling. We conduct upsampling experiments using point-based [43] and transformer-based [38] methods. The results are presented in Tab. 7. It is important to mention that including the HD metric in the evaluation metrics does not effectively reveal the superiority or inferiority of the algorithm in scene-level upsampling experiments since HD only reflects the maximum mismatch between point sets. As can be seen, PU-GCN performs better in 16/32-channel input while PC2-PU is better in 64-channel input. As the density of input points increases, the accuracy of prediction results also improves. This characteristic is more prominent in PC2-PU. The specific visual results are shown in Fig. 5. For the first time, we complete the evaluation of the upsampling task on scene-level point clouds. However, from this result, it can also be seen that compared to traditional 3D object upsampling tasks, the upsampling effect at the scene level is not satisfactory. The details of some large buildings are severely lost, which poses new challenges for the point cloud upsampling task.
Point cloud semantic segmentaion. We utilize pointwise-MLP-based[26], point-convolution-based[53] and voxel-based[11] methods to test our dataset. The results are shown in Tab. 8. It can be seen that point-convolution-based Minkowski U-Net achieve the best results(voxel size is all set to 0.05). As for the pointwise-MLP-based RandLA-Net, the test results for different channel count are not satisfactory. For example, for small categories like traffic signs, there are even cases where the IoU is 0. We speculate that this is because the downsampling strategy used in RandLA-Net is not suitable for segmenting small objects in large scenes, resulting in performance degradation. In point-convolution-based KPConv, compared to the 64-channel data, the 128-channel data did not show a significant improvement. This is because, for the KPConv, with higher-density input point clouds, precise adjustment of the neighborhood radius is essential to prevent performance degradation. The multi-density point clouds we offer facilitate robustness testing of semantic segmentation algorithms across domains, aiding researchers in exploring domain adaptation or generalization studies.
Method Data mIoU normal-$\uparrow$
@@ 5 Multi-task Exploration
In this section, we take upsampling and segmentation as examples to investigate the mutual reinforcement between different tasks. The experimental results show that there is indeed a mutually reinforcing effect between the two subtasks, and that simultaneous training accelerates convergence and improves performance. This also partially demonstrates that our dataset can facilitate 3D multitasking research and represents more possibilities for the future research.
@@ 5.1 The Network
We adopt PU-Dense [2] network architecture as the backbone network and extend a branch for semantic segmentation task. The detailed network architecture is illustrated in Fig. 6. Apart from the original PU-Dense structure, we incorporate transposed convolution [11] to upsample features from different scales of the encoder. These features are then concatenated as input, and the SENet [22] is applied to learn the importance of each channel. Ultimately, a linear classifier is employed to accomplish the semantic segmentation. Besides, we adjust the expansion factor (kernel size) for coordinate generation within the PU-Dense network to accommodate large-scale scenes.
@@ 5.3 Experiments
To explore the relationships between multiple sub-tasks, we adopt the following experimental strategies: training solely with the upsampling branch; training solely with the semantic segmentation branch; and simultaneous training with both the upsampling and semantic segmentation branches. Such a design allows us to quantify the phenomena of inter-task influence, providing a research basis for the design of end-to-end foundation models.
Using data with different point densities from our dataset, we take the coordinate values of the 32-channel point cloud as input, the 128-channel point cloud as the ground truth for upsampling, and the label values of the 32-channel point cloud as the ground truth for semantic segmentation. For a fair comparison, the Adam [30] and SGD [48] optimizer is used with a weight decay of 0.0001 and a default learning rate of 0.002. We use exponentialLR [35] to adjust learning rate and the gamma is 0.5. During the training procedure, random rotation along z-axis is performed for the augmentation. The model is trained for 5 epochs with a batch size of 4 under each setting.
Branch Ups. AR normal-$\uparrow$ Ups. CD normal-$\downarrow$ Seg. mIoU normal-$\uparrow$
@@
The experimental results are shown in Tab. 9. From the experimental results, it can be observed that when the upsampling and segmentation branches are conducted simultaneously, both the Average Recall (AR) of the upsampling task and the mean Intersection over Union (mIoU) of the segmentation task show improvement. They increase from 43.2% to 46.5% and from 43.6% to 45.3%, respectively. Such results demonstrate the phenomenon of mutual promotion among specific downstream sub-tasks. The upsampling task forces the encoder to capture the implicit surface structure of objects, and such information can provide geometric constraints for semantic segmentation. Conversely, the semantic information provided by the semantic segmentation task can also assist the upsampling task in obtaining category-related information, thereby improving the upsampling process.
It is worth noting that the Chamfer Distance (CD) does not meet the expected performance as PU-Dense is a voxel-based method that achieves upsampling by expanding the voxels occupied by the original spatial points. So it does not adequately reflect the quality of the upsampling.
@@
Through simple experimental setups, we can preliminarily leverage the proposed dataset for multi-task learning. This demonstrates the advantage of collecting the dataset from the same scenarios, allowing for the convenient integration of multiple tasks without the need for complex mappings. This also reflects that our dataset can support the exploration of the future end-to-end framework.
6 Discussion and Conclusion
@@
We introduce VEnvision3D, a synthetic perception dataset featuring the exploration of 3D multi-task models, including depth completion, segmentation, upsampling, place recognition, and 3D reconstruction. In a virtual environment, we implement innovative sensor configurations, endowing our dataset with unique characteristics. For instance, it can provide data with varying densities within the same frame, offer surface-sampled data at a city level, and simulate environmental changes. These features aim to contribute to the advancement of the field of point cloud perception. In addition, benefiting from the carefully designed settings in our data collection, we integrat the upsampling task and semantic segmentation task into a multi-task framework. Through experiments, we preliminarily demonstrate the mutually beneficial relationship among different sub-tasks.
However, we also acknowledge the limitations of our dataset. Joint training with only two subtasks can only provide preliminary evidence of the relevance of certain 3D perception tasks. To achieve a truly multi-task end-to-end network, further effective experiments and explorations are needed. We plan to progressively incorporate more sub-tasks in the future to continue our research.
Additionally, although virtual environments possess unique advantages, one might argue that the domain gap between synthetic and real data is a weakness. In our defense, we believe that when synthetic data is used correctly, it can be used to enhance task performance on real data. For example, [14] indicates that the dataset generated by the CARLA simulator can improve the accuracy of point cloud semantic segmentation to a certain extent. [49] showed that augmenting with point clouds generated from Carla simulator can improve bird’s eye view 2D detection performance on the real-world KITTI dataset. [62, 25, 16] developed several domain adaptation methods to mitigate domain shift as much as possible. Furthermore, we also adjusted the sensor settings and added noise on LiDAR to minimize domain shift as much as possible.
In summary, by introducing VEnvision3D, hope that our dataset can aid researchers in conducting more novel preliminary explorations under the current scarcity of 3D datasets, to better realize more grounded concepts in the future. We believe our work is a valuable and necessary first step towards 3D multi-task learning and deserves sharing with the community.